2 As the biologist for a certain crocodile farm, you know that its carrying capacity is 20,000 crocodiles. You initially release 5000 crocodiles into the farm. After 6 weeks, the crocodile population has increased to 7500. In how many weeks will the population reach 10,000? In how many weeks will the population reach 20,000^ prime


1
Expert's answer
2022-05-30T14:57:55-0400
"N(t)=N(0)e^{kt}"

Given "N(0)=5000"


"7500=5000e^{k(6)}"

"6k=\ln 1.5"

"k=\dfrac{1}{6}\ln 1.5"

"N(t)=5000(1.5)^{t\/6}"

a)


"10000=5000(1.5)^{t\/6}"

"(1.5)^{t\/6}=2"

"\dfrac{t}{6}\ln1.5=\ln2"

"t=\dfrac{6\ln 2}{\ln1.5}"

"t=10.257\ weeks"

b)


"20000=5000(1.5)^{t\/6}"

"(1.5)^{t\/6}=4"

"\dfrac{t}{6}\ln1.5=\ln4"

"t=\dfrac{12\ln 2}{\ln1.5}"

"t=20.514\ weeks"